OpenStudy (anonymous):

Find the equation of the line parallel to the line x-2y=-5 containing the point (-3,2)

OpenStudy (paxpolaris):

what's the slope of x-2y=-5 ??

OpenStudy (anonymous):

I think so.. That's all the question gives, so I'm not sure..

OpenStudy (paxpolaris):

no, i mean can you rewrite the equation x-2y=-5 in slope-intercept form (y=mx+b) to give me the slope of that line?

OpenStudy (anonymous):

Is it Y= 1/2x + 5/2?? I'm not sure...

OpenStudy (paxpolaris):

right :) so the slope m would be ...

OpenStudy (anonymous):

1/2?

OpenStudy (paxpolaris):

correct. Parallel lines have the same slope. and you know the parallel line goes through (-3,2) so you can give me the point-slope form for the parallel line: \(y-y_1=m(x-x_1)\)

OpenStudy (anonymous):

y+3=1/2(x-2)?

OpenStudy (paxpolaris):

perfect :D

OpenStudy (anonymous):

what do I do from there?

OpenStudy (anonymous):

I think I figured it out.. Is it Y=1/2x-4?

OpenStudy (paxpolaris):

yes.

OpenStudy (anonymous):

Thank you so much you're a life saver!! I'm sorry I'm probably bothering you, but do you by any chance know how to do another problem to find the perpendicular line with the same slope that contains the points (4,0)?

OpenStudy (paxpolaris):

ok slope for the original line was 1/2. What will be the slope of the perpendicular line?

OpenStudy (paxpolaris):

slope of perpendicular line is the negative reciprocal of the original ...

OpenStudy (anonymous):

The equation of the perpendicular line is still x-2y=-5, so wouldn't the slope still be 1/2? Here's the whole question so I don't confuse you: Find the equation of the line perpendicular to the line x-2y=-5 containing the point (4,0)

OpenStudy (anonymous):

So it would actually be -1/2? since it has to be negative?

OpenStudy (paxpolaris):

x-2y=-5 has slope: 1/2 for a line perpendicular to that the slope has to be -2 (negative reciprocal of 1/2)

OpenStudy (anonymous):

Oh, I see what you're saying... So would the answer be Y=-2x-8, or am I way off?

OpenStudy (paxpolaris):

slope -2 is correct. but when you plug in x=4, you need to get Y=0. Y=-2x-8 doesn't give you that

OpenStudy (anonymous):

Okay, I see what you mean.. I plugged in 4 for x and 0 for y, and got b=8 So would it actually be y=-2x+8?

OpenStudy (paxpolaris):

\[\checkmark\]

OpenStudy (anonymous):

Yay!! I basically can't thank you enough!! I know that we already discussed the first question I had up there, but I was double checking it on my calculator and I think it might actually be y=x+5? Maybe I'm just overthinking it and it really is y=1/2x-4, but..?

OpenStudy (paxpolaris):

y=x+5 passes thru (-3,2) but it is not parallel to x-2y=-5 .... they will intersect. parallel lines MUST have the same slope.

OpenStudy (anonymous):

Oh, yeah... That could be why it wasn't showing up on my calculator, duh they have to have all different points! Thank you so much again!!